# An operator splitting analysis of Wasserstein–Fisher–Rao gradient flows

### v1

@@ Abstract
Wasserstein-Fisher-Rao (WFR) gradient flows have been recently proposed as a powerful sampling tool that combines the advantages of pure Wasserstein (W) and pure Fisher-Rao (FR) gradient flows. Existing algorithmic developments implicitly make use of operator splitting techniques to numerically approximate the WFR partial differential equation, whereby the W flow is evaluated over a given step size and then the FR flow (or vice versa). This works investigates the impact of the order in which the W and FR operator are evaluated and aims to provide a quantitative analysis. Somewhat surprisingly, we show that with a judicious choice of step size and operator ordering, the split scheme can converge to the target distribution faster than the exact WFR flow (in terms of model time). We obtain variational formulae describing the evolution over one time step of both sequential splitting schemes and investigate in which settings the W-FR split should be preferred to the FR-W split. As a step towards this goal we show that the WFR gradient flow preserves log-concavity and obtain the first sharp decay bound for WFR.
@@ 1 Introduction
We consider the task of generating samples from a target probability distribution with density $\pi(x)\propto e^{-V_{\pi}(x)},\enskip x\in\mathbb{R}^{d}$ known up to a normalisation constant. Despite the conceptual simplicity of this task, its efficient implementation when $V_{\pi}$ is multi-modal, the underlying space is high dimensional and/or modes are separated by large distances remains challenging. Given the broad application of sampling to (Bayesian) statistics and statistical machine learning, several avenues have been investigated to derive efficient and scalable algorithms.
@@
A natural way to formulate this task is via gradient flows [ambrosio2008gradient], which can be seen as optimisation of a functional measuring the dissimilarity to $\pi$, typically the Kullback–Leibler (KL) divergence [wibisono2018sampling]. This formulation yields considerable freedom in the design of sampling algorithms; arguably the most well-known being the so-called Wasserstein Gradient flows, hereafter W flows. It is well-known that when $\pi$ satisfies a Log-Sobolev Inequality (LSI), the W flow converges exponentially fast to $\pi$, with rate depending on the Log-Sobolev constant. This highlights an inherent limitation of W flows, namely that when the LSI constant is large (e.g. as is typically the case in multi-modal densities with well separated modes), the convergence rate can be prohibitively slow [schlichting2019poincare].
@@
Recent research efforts have instead considered gradient flows in the Fisher-Rao geometry (FR flow). These are well known in the biological literature as describing the macroscopic properties of a population with varying traits or species, and is sometimes referred to as birth-death or replicator dynamics [kimura_stochastic_1965, schuster1983replicator, Cressman2006]. It is now known that FR flows are intimately connected to mirror descent [chopin2023connection], stochastic filtering [akyildiz2017probabilistic, halder_gradient_2017, Pathiraja2024, DelMoral1997] and Sequential Monte Carlo [us]. The have also recently been exploited to develop sampling algorithms [Nusken2024, maurais2024sampling, Wang2024, chen2023sampling, lu2023birth, Lu2019], most notably due to the fact that it is possible to achieve convergence rates independent of the properties of $V_{\pi}$ [carrillo_fisher-rao_2024, Lu2019]. However, the development of stable and efficient numerical approximations of pure FR flows remains a challenge.
@@
In this work, we consider the so-called Wasserstein–Fisher–Rao (WFR) gradient flow, wherein the metric is given by the direct sum of the Wasserstein and Fisher-Rao metrics. We refer the reader to [liero_optimal_2018] for a rigorous treatment of this metric and its corresponding gradient flow. WFR gradient flows combine the diffusive behaviour of W flows with the birth-death or reactive properties of FR flows and enjoy better convergence properties than both W and FR alone [Lu2019, chen2023sampling]. They have a natural interpretation as combining ‘exploration’ or ‘mutation’ to provide new particles (W flow) with ‘selection’ where particles that are a poor fit to the target are killed [Pathiraja2024]. The convergence properties of WFR flows have been theorised to improve on both W and FR flows, yet the best known results are obtained under strong conditions on the ration between $\pi$ and the initial distribution $\mu_{0}$ [Lu2019, Appendix B] and [lu2023birth, Remark 2.6] and, as we show later, are not sharp. Research into WFR flows is active; several open questions and areas of investigation remain both in terms of algorithmic development and theoretical investigation.
The structure of WFR gradient flows naturally lends itself to numerical methods built on operator splitting [Hundsdorfer2003]. Splitting of the WFR PDE has been considered in several papers when deriving algorithms [us, lu2023birth, Lu2019] but without a precise characterisation of the effect of the splitting choice.
Operator splitting has been previously exploited to build splitting schemes for the kinetic Langevin SDE arising from the W flow to obtain algorithms with no time discretisation bias [chada2023unbiased] and methods with higher accuracy for fixed computational cost [leimkuhler2013rational, bou2010long]. Our work aims to show that splitting schemes can exhibit faster convergence to the target than the exact dynamics and thus a wise choice of splitting can reduce computational cost.
In this work we focus on exact splitting, i.e. we do not consider numerical schemes to approximately solve the W and FR flow and their numerical errors but rather focus on the numerical error introduced by time discretisations via operator splitting. We demonstrate that numerical error introduced by operator splitting can be exploited to accelerate convergence to equilibrium. A particular advantage of this observation is that ordering of the operators does not affect computational cost, and thus speed up can be obtained at no additional cost and without introducing additional numerical approximations.
This is particularly relevant when the computational budget is limited, as the speed up due to splitting is mostly observed in the early iterations. We thus claim that the continuous time process need not be the target of our numerical approximations but rather the splitting scheme which gives us the fastest convergence given $\pi,\mu_{0}$. This then paves the way for the development of practical numerical algorithms (i.e. space-time discretisations) that approximate the operator split WFR flow rather than the fully coupled WFR gradient flow. There are a myriad schemes for numerically approximating both W [NEURIPS2019_6a8018b3_salim, salim2020wasserstein, mou2021high] and FR [nusken2024transport, maurais2024sampling, Wang2024, korba2022adaptive, chen2023sampling] flows and we leave their detailed investigation to future work.
@@
  • We derive variational formulae (PDEs) that quantify the impact of operator splitting as function of the step size (Section 2) and demonstrate that a carefully chosen operator splitting method can yield theoretical convergence rates that are faster than the fully continuous time WFR flow (Section 6) without increasing the computational cost.
  • To do so, we obtain the first convergence result for the continuous time WFR flow which clearly shows that the rate of convergence is the sum of the rate of the W flow and that of the FR flow (Section 5). This result is obtained after establishing that the WFR preserves log-concavity thanks to the strong regularising properties of the FR flow (Section 4).
  • We provide as motivation a precise quantitative characterisation in the multivariate Gaussian case (Section 3) which allows us to gain some intuition on general guidelines on the choice of initial distribution $\mu_{0}$ to obtain faster convergence (Section 6).
@@ 2 Operator Splitting for the WFR PDE
The gradient flow of $\mathrm{KL}(\mu||\pi)$ w.r.t. the geometry induced by the Wasserstein–Fisher–Rao distance is given by the following PDE [Lu2019, Theorem 3.1]
@@
and $f_{\textrm{W}}$ and $f_{\textrm{FR}}$ correspond to the Wasserstein and Fisher–Rao operators respectively, and takes the form of a reaction-diffusion equation. This PDE is a combination of the Wasserstein (W) flow and of the Fisher–Rao (FR) flow and enjoys better convergence properties than both W and FR alone [Lu2019, chen2023sampling].
The WFR PDE (2.1) corresponds to the underlying continuous time dynamics of many recent sampling methods as highlighted in [us]. Recent methods involve using a variety of numerical schemes e.g. interacting particle based methods [Lu2019, lu2023birth, chen2023sampling] and kernel based methods [maurais2024sampling, Nusken2024, Wang2024]. A well-known approach to discretising PDEs involving sums of operators is operator splitting [Hundsdorfer2003, blanes2024splitting] which has been applied to the W flow to split advection and diffusion and led to a vast number of numerical approximations of Langevin dynamics [abdulle2015long, duncan2017nonreversible, chada2023unbiased, leimkuhler2013rational, bou2010long]. Such operator splitting is particularly natural for the WFR PDE, where the W flow is characterised by advection and diffusion operators whilst the FR flow is characterised by reaction dynamics. Splitting of the WFR PDE has been considered in several papers when deriving algorithms [us, lu2023birth, Lu2019] but without a precise characterisation of the effect of the splitting choice.
Throughout this manuscript, we focus on time discretisation of (2.1) via operator splitting under the assumption that the operators are evaluated exactly over a single time step, i.e. without further discretisations using either determinstic or Monte Carlo type approximations. Our motivation for doing so is to quantify biases due solely to splitting error, thereby giving rise to a new PDE which can potentially yield faster (continuous time) convergence rates than the original WFR PDE. This opens the door for the development and study of numerical schemes whose underlying dynamics correspond to the split WFR flow rather than its fully continuous time counterpart. In particular, we assume that the solution operator of the W flow for a time step $\gamma$, $S_{\textrm{W}}(\gamma,v)(x)$, can be evaluated exactly. Similarly, we assume the solution operator of the FR flow given by (e.g. [chen2023sampling, Appendix B.1])
@@
can also be evaluated exactly. A commonly used first order accurate splitting method is sequential splitting or Lie-Trotter splitting (see, e.g., [blanes2024splitting] for a recent review), where each operator is solved over a step size in an alternating manner. More specifically, consider a time-discretisation $t_{0}=0<t_{1}<t_{2}\cdots<t_{M}=T$ with $t_{n+1}-t_{n}=\gamma\enskip\forall\enskip i=1,2,\dots M$. A sequential splitting applied to (2.1) takes the form
$\hat{\nu}_{1}(x;\gamma)$ $=S_{\textrm{W}}(\gamma,\mu_{0})(x)$ (2.6)
$\nu_{1}(x;\gamma)$ $=S_{\textrm{FR}}(\gamma,\hat{\nu}_{1})(x)$
$\hat{\nu}_{2}(x;\gamma)$ $=S_{\textrm{W}}(\gamma,\nu_{1})(x)$
$\nu_{2}(x;\gamma)$ $=S_{\textrm{FR}}(\gamma,\hat{\nu}_{2})(x)$
$\vdots$
where $S_{\textrm{FR}}(\gamma,v)$ and $S_{\textrm{W}}(\gamma,v)$ denote the solution operators corresponding to $f_{\textrm{FR}}(\mu)$ and $f_{\textrm{W}}(\mu)$ acting on $v$ over time interval of size $\gamma$, respectively. We refer to (2.6) as Wasserstein then Fisher–Rao, or W-FR for short. It is known that the $n$th iterate of the splitting scheme (2.6), $\nu_{n}$, converges to the solution of (2.1) at time $t=n\gamma$, $\mu_{n\gamma}$, as $\gamma\rightarrow 0$ (at rate $\mathcal{O}(\gamma)$) under mild conditions on $f_{\textrm{FR}}$, $f_{\textrm{W}}$ and $\pi$ (see e.g. [Hundsdorfer2003]).
@@
$\hat{\eta}_{1}(x;\gamma)$ $=S_{\textrm{FR}}(\gamma,\mu_{0})(x)$ (2.7)
$\eta_{1}(x;\gamma)$ $=S_{\textrm{W}}(\gamma,\hat{\eta}_{1})(x)$
$\hat{\eta}_{2}(x;\gamma)$ $=S_{\textrm{FR}}(\gamma,\eta_{1})(x)$
$\eta_{2}(x;\gamma)$ $=S_{\textrm{W}}(\gamma,\hat{\eta}_{2})(x)$
$\vdots$
We refer to (2.7) as Fisher–Rao then Wasserstein, or FR-W for short. When the operators $S_{\textrm{FR}}$ and $S_{\textrm{W}}$ do not commute, it is known that $\eta_{i}(x;\gamma)\neq\nu_{i}(x;\gamma)$ for $\gamma>0$ and $i=1,2,3,\dots$ [Hundsdorfer2003]. It is expected that for most target distributions, $S_{\textrm{FR}}$ and $S_{\textrm{W}}$ will not commute; the following example demonstrates this occurs even in the multivariate Gaussian case.
In the remainder of the manuscript, we rely on the following assumptions on the target distribution $\pi$ and the distribution at which the WFR PDE (2.1) is initialised $\mu_{0}$ in developing analytic results.
@@
Remark 2.1.
One could also apply a splitting scheme for pure W to split the advection and diffusion parts (leading to proximal samplers, see e.g. [wibisono2018sampling, Section 4.2]). Similarly, one could also apply a splitting scheme to the pure FR flow by identifying an exploration step and an exploitation step as in [chen_efficient_2024, Section 4]. We leave analysis of these alternative splitting schemes for future work.
@@ 2.1 Variational Formula: W-FR
We begin by analysing the evolution obtained by using the splitting scheme in (2.6). In particular, we obtain a PDE which describes the evolution of the splitting solution over time interval $0,\gamma]$. The advantage of this characterisation is that we can identify a perturbation to the exact dynamics and more easily characterise how splitting biases modify convergence speed. The traditional route to analysing decay of $\mathrm{KL}$ of the WFR dynamics in e.g. [us, Proposition 3.1], [lu2023birth, Remark 2.6], [Lu2019, Appendix B] is not suited to the splitting schemes as it does not allow us to distinguish the order in which the W and FR solution operators are applied.
The following proposition quantifies the variation of $\nu_{\gamma}$ with respect to the step size $\gamma$ using standard variational calculus tools. For the sake of notational brevity, we rely on the shorthand notation $\nu_{\gamma}$ to refer to $\nu_{1}(x;\gamma)$.
@@
Sequential split (W-FR) PDE: Let us denote $\nu_{\gamma}=\nu_{1}(x;\gamma)$ and let Assumption 1 hold. The variation of one sequential split step in the order W-FR (2.6) of size $\gamma$ corresponds to the PDE
$\partial_{\gamma}\nu_{\gamma}=f_{\textrm{W}}(\nu_{\gamma})+f_{\textrm{FR}}(\nu_{\gamma})+(e^{\gamma}-1)\nu_{\gamma}(|\nabla g(\nu_{\gamma})|^{2}-\mathbb{E}_{\nu_{\gamma}}[|\nabla g|^{2}]),$ (2.9)
@@
Proof.
Here we use the notation $\rho_{\gamma}$ to refer to the unnormalised form of $\nu_{\gamma}$. Denote by $f_{\textrm{G}}$ the unnormalised FR operator,
$f_{\textrm{G}}(\mu):=\mu\log\frac{\pi}{\mu}.$
$\rho_{\gamma}=S_{\textrm{G}}(\gamma,S_{\textrm{W}}(\gamma,\mu_{0})),$
where $S_{\textrm{G}}(\gamma,v)$ denotes the abstract solution operator of $\partial_{\gamma}\rho_{\gamma}=f_{\textrm{G}}(\rho_{\gamma})$ initialised at $v$ after time $\gamma$ and $S_{\textrm{W}}(\gamma,v)$ denotes the solution operator of $\partial_{\gamma}\rho_{\gamma}=f_{\textrm{W}}(\rho_{\gamma})$. Using the chain rule, the first variation of $\rho_{\gamma}$ with respect to $\gamma$ is (see also (3.4) in [descombes_lietrotter_2013])
$\partial_{\gamma}\rho_{\gamma}=f_{\textrm{G}}(\rho_{\gamma})+\partial_{2}S_{\textrm{G}}(\gamma,v_{\gamma})f_{\textrm{W}}(v_{\gamma}),$ (2.10)
where $v_{\gamma}:=S_{\textrm{W}}(\gamma,\mu_{0})$ and $\partial_{i}$ denotes the directional derivative w.r.t. the $i$-th argument. Due to [chen2023sampling, Appendix B.1], the solution operator $S_{\textrm{G}}$ has an explicit form given by $S_{\textrm{G}}(\gamma,v)=\pi^{1-e^{-\gamma}}v^{e^{-\gamma}}$, so that $\partial_{2}S_{\textrm{G}}(\gamma,v)h$, i.e. the directional derivative w.r.t. the second argument $v\in\mathcal{P}_{2}^{ac}$ in the direction $h\in\mathcal{T}$ is easily computed as,
$\partial_{2}S_{\textrm{G}}(\gamma,v)h=\left.\frac{d}{d\epsilon}S_{\textrm{G}}(\gamma,v+\epsilon h)\right|_{\epsilon=0}$ $=\left.\frac{d}{d\epsilon}\pi^{1-e^{-\gamma}}(v+\epsilon h)^{e^{-\gamma}}\right|_{\epsilon=0}$
$=e^{-\gamma}\pi^{1-e^{-\gamma}}v^{e^{-\gamma}-1}h$
$=e^{-t}S_{\textrm{G}}(t,v)hv^{-1}.$
It is easy to see that $\partial_{2}S_{\textrm{G}}(\gamma,v)h$ is a linear operator and that $x\mapsto\partial_{2}S_{\textrm{G}}(\gamma,v)(x)=e^{-t}(\pi(x)/v(x))^{1-e^{-t}}$ is continuous if $\pi,v$ are continuous and $v(x)>0$. Substituting back into (2.10) yields
$\partial_{\gamma}\rho_{\gamma}=f_{\textrm{G}}(\rho_{\gamma})+e^{-\gamma}S_{\textrm{G}}(\gamma,v_{\gamma})f_{\textrm{W}}(v_{\gamma})v_{\gamma}^{-1}.$ (2.11)
Notice that the r.h.s can be written entirely in terms of $\rho_{\gamma}$ since $S_{\textrm{G}}$ is an invertible operator, with inverse given by
Since $\rho_{\gamma}=S_{\textrm{G}}(\gamma,v_{\gamma})$ where $v_{\gamma}=S_{\textrm{W}}(\gamma,\mu_{0})$, we have using the above that,
$v_{\gamma}=\pi^{1-e^{\gamma}}\rho_{\gamma}^{e^{\gamma}}.$
Then $f_{\textrm{W}}(v_{\gamma})$ can be expressed in terms of $\rho_{\gamma}$ since, using the shorthand notation $w:=\nabla\log\frac{\rho_{\gamma}}{\pi}$,
$f_{\textrm{W}}(v_{\gamma})$ $=\nabla\cdot\left(v_{\gamma}\nabla\log\frac{v_{\gamma}}{\pi}\right)$
$=e^{\gamma}\nabla\cdot\left(\pi^{-e^{\gamma}+1}\rho_{\gamma}^{e^{\gamma}}w\right)$
$=e^{\gamma}(\pi^{-e^{\gamma}+1}\nabla\cdot\left(\rho_{\gamma}^{e^{\gamma}-1}\rho_{\gamma}w\right)+\rho_{\gamma}^{e^{\gamma}}w\cdot\nabla(\pi^{-e^{\gamma}+1}))$
$=e^{\gamma}(\pi^{-e^{\gamma}+1}\rho_{\gamma}^{e^{\gamma}-1}\nabla\cdot\left(\rho_{\gamma}w\right)+\pi^{-e^{\gamma}+1}\rho_{\gamma}w\cdot\nabla(\rho_{\gamma}^{e^{\gamma}-1})+\rho_{\gamma}^{e^{\gamma}}w\cdot\nabla(\pi^{-e^{\gamma}+1}))$
$=e^{\gamma}\pi^{-e^{\gamma}+1}\rho_{\gamma}^{e^{\gamma}-1}(f_{\textrm{W}}(\rho_{\gamma})+\underbrace{\rho_{\gamma}^{2-e^{\gamma}}w\cdot\nabla(\rho_{\gamma}^{e^{\gamma}-1})+\rho_{\gamma}\pi^{e^{\gamma}-1}w\cdot\nabla(\pi^{-e^{\gamma}+1})}_{=:R}),$
and notice that $R$ can be simplified further as
$R$ $=\rho_{\gamma}\left(\rho_{\gamma}^{1-e^{\gamma}}w\cdot\nabla(\rho_{\gamma}^{e^{\gamma}-1})+\pi^{e^{\gamma}-1}w\cdot\nabla(\pi^{-e^{\gamma}+1})\right)$
$=\rho_{\gamma}\left((e^{\gamma}-1)\rho_{\gamma}^{-1}w\cdot\nabla\rho_{\gamma}-w\cdot(e^{\gamma}-1)\pi^{-1}\nabla\pi+\nabla\cdot w\right)$
$=(e^{\gamma}-1)\rho_{\gamma}|w|^{2},$
so that
$f_{\textrm{W}}(v_{\gamma})=e^{\gamma}\pi^{-e^{\gamma}+1}\rho_{\gamma}^{e^{\gamma}-1}(f_{\textrm{W}}(\rho_{\gamma})+(e^{\gamma}-1)\rho_{\gamma}|w|^{2}).$
Substituting all the derived expressions back into (2.11) yields
$\partial_{\gamma}\rho_{\gamma}$ $=f_{\textrm{G}}(\rho_{\gamma})+e^{-\gamma}e^{\gamma}\pi^{-e^{\gamma}+1}\rho_{\gamma}^{e^{\gamma}-1}(f_{\textrm{W}}(\rho_{\gamma})+(e^{\gamma}-1)\rho_{\gamma}|w|^{2})\pi^{e^{\gamma}-1}\cdot\rho_{\gamma}^{1-e^{\gamma}}$
$=f_{\textrm{W}}(\rho_{\gamma})+f_{\textrm{G}}(\rho_{\gamma})+(e^{\gamma}-1)\rho_{\gamma}|w|^{2}.$
Then by an application of [us, Lemma A.1] to obtain the PDE for the normalised density $\nu_{\gamma}$, the claim holds.
The above proposition shows that splitting introduces a perturbation to the WFR PDE (2.1) of the form
$f_{\textrm{P}}(\nu_{\gamma}):=\nu_{\gamma}\left(\left|\nabla g(\nu_{\gamma})\right|^{2}-\mathbb{E}_{\nu_{\gamma}}\left[\left|\nabla g(\nu_{\gamma})\right|^{2}\right]\right),$ (2.12)
where $g$ is as defined in (2.4), whose contribution is modulated by an additional factor $(e^{\gamma}-1)$ which increases monotonically with step size $\gamma$. Somewhat remarkably, a simple splitting procedure is sufficient to generate an acceleration under certain conditions as we will show in Sections 3 and 6.

### v2

# arXiv:2511.18060v2[stat.ML] 27 Feb 2026

@@ Abstract
Wasserstein-Fisher-Rao (WFR) gradient flows have been recently proposed as a powerful sampling tool that combines the advantages of pure Wasserstein (W) and pure Fisher-Rao (FR) gradient flows. Existing algorithmic developments implicitly make use of operator splitting techniques to numerically approximate the WFR partial differential equation, whereby the W flow is evaluated over a given step size and then the FR flow (or vice versa). This works investigates the impact of the order in which the W and FR operator are evaluated and aims to provide a quantitative analysis. Somewhat surprisingly, we show that with a judicious choice of step size and operator ordering, the split scheme can converge to the target distribution faster than the exact WFR flow (in terms of model time). We obtain variational formulae describing the evolution over one time step of both splitting schemes and investigate in which settings the W-FR split should be preferred to the FR-W split. As a step towards this goal we show that the WFR gradient flow preserves log-concavity and obtain the first sharp decay bound for WFR flow.
@@ 1 Introduction
We consider the task of generating samples from a target probability distribution known up to a normalisation constant with density $\pi(x)\propto e^{-V_{\pi}(x)},\enskip x\in\mathbb{R}^{d}$. Despite the conceptual simplicity of this task, its efficient implementation when $V_{\pi}$ is multi-modal, the underlying space is high dimensional and/or modes are separated by large distances remains challenging. Given the broad application of sampling to (Bayesian) statistics and statistical machine learning, several avenues have been investigated to derive efficient and scalable algorithms.
@@
A natural way to formulate this task is via gradient flows, which can be seen as optimisation of a functional measuring the dissimilarity to $\pi$, typically the Kullback–Leibler (KL) divergence [WIB18]. This formulation yields considerable freedom in the design of sampling algorithms; arguably the most well-known being the so-called Wasserstein Gradient flow, hereafter W flow. It is well-known that when $\pi$ satisfies a Log-Sobolev Inequality (LSI), the W flow converges exponentially fast to $\pi$, with rate depending on the Log-Sobolev constant. This highlights an inherent limitation of W flows, namely that when the LSI constant is large (e.g. as is typically the case in multi-modal densities with well separated modes), the convergence rate can be prohibitively slow [SCH19].
@@
Recent research efforts have instead considered gradient flows in the Fisher–Rao geometry (FR flow). These are well known in the biological literature as describing the macroscopic properties of a population with varying traits or species, and is sometimes referred to as birth-death or replicator dynamics [KIM65, SS83, CHR06]. It is known that FR flows are intimately connected to mirror descent [CCK24], stochastic filtering [AKY17, HG17, PW24, DEL97] and sequential Monte Carlo [CP25]. They have also recently been exploited to develop sampling algorithms [NÜS24, MM24, WN24, CHH+23, LSW23, LLN19], most notably due to the fact that it is possible to achieve convergence rates independent of the properties of $V_{\pi}$ [CCH+24, LLN19]. However, the development of stable and efficient numerical approximations of pure FR flows remains a challenge.
@@
In this work, we consider the so-called Wasserstein–Fisher–Rao (WFR) gradient flow, wherein the metric is given by the direct sum of the Wasserstein and Fisher-Rao metrics. We refer the reader to [LMS18] for a rigorous treatment of this metric and its corresponding gradient flow. WFR gradient flows combine the diffusive behaviour of W flows with the birth-death or reactive properties of FR flows and enjoy better convergence properties than both W and FR alone [LLN19, CHH+23]. They have a natural interpretation as combining ‘exploration’ or ‘mutation’ to provide new particles (W flow) with ‘selection’ where particles that are a poor fit to the target are killed [PW24]. The convergence properties of WFR flows have been theorised to improve on both W and FR flows, yet the best known results are obtained under strong conditions on the ratio between $\pi$ and the initial distribution $\mu_{0}$ [LLN19, Appendix B] and [LSW23, Remark 2.6] and, as we show later, are not sharp. Research into WFR flows is active; several open questions and areas of investigation remain both in terms of algorithmic development and theoretical investigation.
The WFR partial differential equation (PDE) corresponds to the underlying continuous time dynamics of many recent sampling methods as highlighted in [CP25]. A well-known approach to discretising PDEs involving sums of operators is operator splitting [HV03, BCM24], whereby each operator is solved separately in sequence over a chosen time horizon (referred to as a step size herein). Such approaches have been applied to the W flow to split advection and diffusion and led to a vast number of numerical approximations of Langevin dynamics [AVZ15, DPZ17, CLP+23, LM13, BO10]. Operator splitting has also been previously exploited to build splitting schemes for the kinetic Langevin SDE to obtain algorithms with no time discretisation bias [CLP+23] and methods with higher accuracy for fixed computational cost [LM13, BO10]. Such operator splitting is particularly natural for the WFR PDE, where the W flow is characterised by advection and diffusion operators whilst the FR flow is characterised by reaction dynamics. Splitting of the WFR PDE has been considered in several papers when deriving algorithms [CP25, LSW23, LLN19] but without a precise characterisation of the effect of the splitting choice.
Here we demonstrate that in the context of PDE based gradient flows, a carefully chosen splitting can exhibit faster convergence to the target than the exact dynamics. In particular, we demonstrate that numerical error introduced by operator splitting can be exploited to accelerate convergence to equilibrium. A particular advantage of this observation is that ordering of the operators does not affect computational cost, and thus speed-up can be obtained without introducing additional numerical approximations. This is particularly relevant when the computational budget is limited, as the speed-up due to splitting is mostly observed in the early iterations. We thus claim that the continuous time process need not be the target of algorithmic implementations, but rather the splitting scheme and associated step size/time horizon over which each operator is solved which yields the fastest convergence given $\pi,\mu_{0}$. Alternatively, existing algorithmic implementations that involve alternating between a single numerical discretisation step of W and FR operators may benefit from instead having a sequence of $n$ W steps, followed by a sequence of $n$ FR steps (or vice versa). In this work, we focus on exact splitting, i.e. we do not consider numerical schemes to approximately solve the W and FR flow. There are a myriad schemes for numerically approximating both W [SKR19, SKL20, MMW+21] and FR [NVP+24, MM24, WN24, KP22, CHH+23] flows, we leave their detailed investigation to future work.
@@
  • We derive variational formulae (PDEs) that quantify the impact of operator splitting as function of the step size (Section 2) and demonstrate that a carefully chosen operator splitting method (i.e. ordering of operators) can yield upper bounds on convergence rates that are faster than the continuous time WFR flow (Sections  3 &  5.2) without increasing computational cost.
  • In doing so, we obtain the first convergence result for the continuous time WFR flow which shows that the rate of convergence is the sum of the rate of the W flow and that of the FR flow (Section 5.1), as previously conjectured in [DP23]. This result is obtained for the case of strongly log-concave targets, after establishing that the WFR preserves log-concavity thanks to the strong regularising properties of the FR flow (Section 4).
  • We provide as motivation a precise quantitative characterisation of the effects of operator ordering in the multivariate Gaussian case (Section 3).
@@ 2 Operator Splitting for the WFR PDE
The gradient flow of $\mathrm{KL}(\mu||\pi)$ w.r.t. the geometry induced by the Wasserstein–Fisher–Rao distance is given by the following PDE [LLN19, Theorem 3.1]
@@
and $f_{\textrm{W}}$ and $f_{\textrm{FR}}$ correspond to the Wasserstein and Fisher–Rao operators respectively, and takes the form of a reaction-diffusion equation. This PDE is a combination of the Wasserstein (W) flow and of the Fisher–Rao (FR) flow and enjoys better convergence properties than both W and FR alone [LLN19, CHH+23].
Throughout this manuscript, we focus on time discretisation of (2.1) via operator splitting under the assumption that the operators are evaluated exactly over a single time step, i.e. without further discretisations using either determinstic or Monte Carlo type approximations. Our motivation for doing so is to quantify biases due solely to splitting error, thereby giving rise to a new PDE which can potentially yield faster (continuous time) convergence rates than the original WFR PDE. This opens the door for the development and study of numerical schemes whose underlying dynamics correspond to the split WFR flow rather than its fully continuous time counterpart. In particular, we assume that the solution operator of the W flow over a time step $\gamma$, $S_{\textrm{W}}(\gamma,v)(x)$, can be evaluated exactly. Similarly, we assume the solution operator of the FR flow given by (e.g. [CHH+23, App. B.1])
@@
can also be evaluated exactly. A commonly used first order accurate splitting method is sequential splitting or Lie-Trotter splitting (see, e.g., [BCM24] for a recent review), where each operator is solved over a step size in an alternating manner. More specifically, consider a time-discretisation $t_{0}=0<t_{1}<t_{2}\cdots<t_{M}=T$ with $t_{n+1}-t_{n}=\gamma\enskip\forall\enskip i=1,2,\dots M$. A sequential splitting applied to (2.1) takes the form
$\hat{\nu}_{i}(x;\gamma)$ $=S_{\textrm{W}}(\gamma,\nu_{i-1})(x)\rightarrow\nu_{i}(x;\gamma)=S_{\textrm{FR}}(\gamma,\hat{\nu}_{i})(x),\quad i=2,3,\dots$ (2.6)
with $\nu_{0}=\mu_{0}$ and where $S_{\textrm{FR}}(\gamma,v)$ and $S_{\textrm{W}}(\gamma,v)$ denote the solution operators corresponding to $f_{\textrm{FR}}$ and $f_{\textrm{W}}$ respectively, acting on $v$ over time interval of size $\gamma$. We refer to (2.6) as Wasserstein then Fisher–Rao scheme, or W-FR for short. We note that the step size $\gamma$ for the W and FR components need not be the same. It is known that
$\nu_{n}\rightarrow\mu_{n\gamma},\quad\gamma\rightarrow 0$
at rate $\mathcal{O}(\gamma)$ under mild conditions on $f_{\textrm{FR}}$, $f_{\textrm{W}}$ and $\pi$ (e.g. [HV03]).
@@
$\hat{\eta}_{i}(x;\gamma)$ $=S_{\textrm{FR}}(\gamma,\eta_{i-1})(x)\rightarrow\eta_{i}(x;\gamma)=S_{\textrm{W}}(\gamma,\hat{\eta}_{i})(x),\quad i=2,3,\dots$ (2.7)
We refer to (2.7) as the Fisher–Rao then Wasserstein scheme, or FR-W for short. When the operators $S_{\textrm{FR}}$ and $S_{\textrm{W}}$ do not commute, it is known that $\eta_{i}(x;\gamma)\neq\nu_{i}(x;\gamma)$ for $\gamma>0$ and $i=1,2,3,\dots$ [HV03]. It is expected that for most target distributions, $S_{\textrm{FR}}$ and $S_{\textrm{W}}$ will not commute; indeed this can be straightforwardly verified even in the multivariate Gaussian case.
Remark 2.1.
One could also apply a splitting scheme for pure W to split the advection and diffusion parts (leading to proximal samplers, see e.g. [WIB18, Section 4.2]). Similarly, one could also apply a splitting scheme to the pure FR flow by identifying an exploration step and an exploitation step as in [CHH+24, Section 4]. We leave analysis of these alternative splitting schemes for future work.
Throughout we rely on the following assumptions on the target distribution $\pi$ and the distribution at which the WFR PDE (2.1) is initialised $\mu_{0}$ in developing analytic results.
@@
In the remainder of the manuscript, we use the notation $\nu_{i}(x)$ and $\eta_{i}(x)$ in place of $\nu_{i}(x;\gamma)$ and $\eta_{i}(x;\gamma)$ respectively as we assume a step size of $\gamma$ throughout. In the next section, we investigate the behaviour of these quantities as $\gamma$ varies.
@@ 2.1 Variational Formula: W-FR
We begin by analysing the splitting scheme in (2.6). In particular, we obtain a PDE which describes the evolution of the splitting solution over time interval $0,\gamma]$. This can equivalently be seen as describing how $\nu_{1}$ varies with respect to the step size. The advantage of this characterisation is that we can identify a perturbation to the exact dynamics and more easily characterise how splitting biases modify convergence speed. The traditional route to analysing decay of $\mathrm{KL}$ of the WFR dynamics in e.g. [CP25, Proposition 3.1], [LSW23, Remark 2.6], [LLN19, Appendix B] is not suited to the splitting schemes as it does not allow us to distinguish the order in which the W and FR solution operators are applied.
The following proposition quantifies the variation of $\nu_{\gamma}$ with respect to the step size $\gamma$ using standard variational calculus tools. Its proof can be found in Appendix A.1.
@@
Sequential split (W-FR) PDE: Let us denote $\nu_{1}=\nu_{1}(x;\gamma)$ and let Assumption 1 hold. The variation of one sequential split step in the order W-FR (2.6) of size $\gamma$ corresponds to the PDE
$\partial_{\gamma}\nu_{1}$ $=f_{\textrm{W}}(\nu_{1})+f_{\textrm{FR}}(\nu_{1})+(e^{\gamma}-1)f_{\textrm{P}}(\nu_{1})$ (2.9)
$f_{\textrm{P}}(\nu_{1})$ $:=\nu_{1}\left(\left|\nabla g(\nu_{1})\right|^{2}-\mathbb{E}_{\nu_{1}}\left[\left|\nabla g(\nu_{1})\right|^{2}\right]\right),$ (2.10)
@@
The above proposition shows that splitting introduces a perturbation of Fisher-Rao structure to the WFR PDE, whose contribution is modulated by an additional factor $(e^{\gamma}-1)$ which increases monotonically with step size $\gamma$. Somewhat remarkably, this simple splitting procedure is sufficient to generate an acceleration under certain conditions, as we will demonstrate in Section 3 for the Gaussian case and in Section 5 for the log-concave case.